# Video Diffusion Transformers are In-Context Learners

### v2

@@ Abstract
This paper investigates a solution for enabling in-context capabilities ofmain video diffusion transformers, with minimal tuning required for activation. Specifically, we propose a simple pipeline to leverage in-context generation: (i) concatenate videos along spacial or time dimension, (ii) jointly caption multi-scene video clips from one source, and (iii) apply task-specific fine-tuning using carefully curated small datasets. Through a series of diverse controllable tasks, we demonstrate qualitatively that existing advanced text-to-video models can effectively perform in-context generation. Notably, it allows for the creation of consistent multi-scene videos exceeding 30 seconds in duration, without additional computational overhead. Importantly, this method requires no modifications to the original models, results in high-fidelity video outputs that better align with prompt specifications and maintain role consistency. Our framework presents a valuable tool for the research community and offers critical insights for advancing product-level controllable video generation systems. The data, code, and model weights are publicly available at: https://github.com/feizc/Video-In-Context.
@@ 2.2 Diffusion Transformers
Transformer architecture [82] has garnered significant attention for its success in language modeling [65, 66, 69] and substantial potential across a wide range of computer vision applications, including image classification [11, 36, 80], object detection [53, 88, 89, 6], and semantic segmentation [109, 92, 78], among others [54, 36, 21, 17, 18, 16]. Building on this success, diffusion Transformers [60] and their variants [1, 24, 26, 27, 25] have replaced the traditional convolutional U-Net backbone [72] with Transformer-based architectures. This shift has resulted in improved scalability and simpler parameter expansion compared to U-Net-based diffusion models. Such advantages in scalability have been particularly evident in areas like video generation [56], image generation [7], speech generation [52], and video generation [98]. Furthermore, recent innovations, such as StableAudio [14] and FluxMusic [23, 22, 20], have explored the application of the DiT architecture in the generation of audio and music.

### v3

@@ Abstract
This paper investigates a solution for enabling in-context capabilities ofmain video diffusion transformers, with minimal tuning required for activation. Specifically, we propose a simple pipeline to leverage in-context generation: (i) concatenate videos along spacial or time dimension, (ii) jointly caption multi-scene video clips from one source, and (iii) apply task-specific fine-tuning using carefully curated small datasets. Through a series of diverse controllable tasks, we demonstrate qualitatively that existing advanced text-to-video models can effectively perform in-context generation. Notably, it allows for the creation of consistent multi-scene videos exceeding 30 seconds in duration, without additional computational overhead. Importantly, this method requires no modifications to the original models, results in high-fidelity video outputs that better align with prompt specifications and maintain role consistency. Our framework presents a valuable tool for the research community and offers critical insights for advancing product-level controllable video generation systems. The code and model weights are publicly available.
@@ 2.2 Diffusion Transformers
Transformer architecture [85] has garnered significant attention for its success in language modeling [69, 70, 72] and substantial potential across a wide range of computer vision applications, including image classification [11, 38, 83], object detection [55, 91, 92, 6], and semantic segmentation [112, 95, 81], among others [56, 38, 27, 17, 29, 16, 26]. Building on this success, diffusion Transformers [63] and their variants [1, 22, 24, 25, 23] have replaced the traditional convolutional U-Net backbone [75] with Transformer-based architectures. This shift has resulted in improved scalability and simpler parameter expansion compared to U-Net-based diffusion models. Such advantages in scalability have been particularly evident in areas like video generation [58], image generation [7], speech generation [54], and video generation [101, 28]. Furthermore, recent innovations, such as StableAudio [14] and FluxMusic [21, 20, 18, 67], have explored the application of the DiT architecture in the generation of audio and music.